OpenStudy (anonymous):

Need help with a question asap! 1. Which ordered pair is a solution of the equation y=x-2? a. (3,1) b. (1,3) c. (-1,3) d. (3,-1)

OpenStudy (texaschic101):

this is just a matter of subbing to see which one is true. Your points are always set up like this : (x,y) y = x - 2 (a.) (3,1)...x = 3 and y = 1 1 = 3 - 2 1 = 1 (correct)......this is is a solution